# ASTGI: Adaptive Spatio-Temporal Graph Interactions for Irregular Multivariate Time Series Forecasting

### v1

@@ 1 Introduction
In the field of time series analysis, researchers have developed various mature forecasting models for regular multivariate time series(Qiu et al., 2024; Wu et al., 2024; Qiu et al., 2025c; Sun et al., 2024, 2025a; Niu et al., 2025), which are characterized by synchronized observations sampled at a fixed frequency (Wu et al., 2025a; Qiu et al., 2025d; Wu et al., 2025b; Qiu et al., 2025a; Liu et al., 2025a; Qiu et al., 2025e). However, in many critical scientific and engineering domains, time series data often exhibit irregular characteristics, posing unique forecasting challenges. Irregular Multivariate Time Series Forecasting (IMTSF) is a core problem in these areas, with wide-ranging applications from monitoring vital signs in intensive care units to tracking the evolution of environmental indicators in climate science. This type of data is fundamentally defined by two characteristics: 1) intra-series irregularity, where observations of the same variable occur at unequal time intervals, and 2) inter-series asynchrony, where observation timestamps are misaligned across different variables. These traits make it challenging to directly apply existing models designed for regular time series (Dai et al., 2024a; Qiu et al., 2025b; Hu et al., 2025; Liu et al., 2025b; Dai et al., 2024b; Sun et al., 2025b). Furthermore, despite existing research on irregular time series prediction, current methods are commonly limited by two core challenges, which impede their potential application in complex real-world scenarios.
@@
The first challenge is how to accurately represent the original information of an irregular time series without introducing information distortion. An accurate representation must preserve the original sampling pattern, as this pattern itself contains key information about the system’s dynamics (Shukla & Marlin, 2020; Li & Marlin, 2020). Conversely, any artificial alteration to the original data structure introduces information distortion, corrupting this intrinsic information and ultimately compromising the model’s predictive performance (Ansari et al., 2023). There are three main representation paradigms, each with its own limitations. 1) Interpolation-based Methods (Figure 1b): This paradigm transforms an irregular series into an equally spaced series through numerical interpolation (Shukla & Marlin, 2021). However, this approach generates artificial data points that were not actually observed, which can introduce bias and distort the original sampling distribution (Shukla & Marlin, 2019; Zhang et al., 2023). 2) Time-aligned-based Methods (Figure 1c): This paradigm maps the observations of all variables to a unified timeline and fills in the missing values (Brouwer et al., 2019). Its main drawback is the loss of precise information about the time intervals between the original observations (Che et al., 2016; Rubanova et al., 2019). 3) Patch-aligned-based Methods (Figure 1d): This paradigm divides the time series into multiple patches (Nie et al., 2023). However, a rigid division may disrupt the continuity of information. Furthermore, intra-patch aggregation can smooth out critical, fine-grained dynamics (Zhang et al., 2024; Luo et al., 2025; Liu et al., 2025c).
@@
The second challenge lies in how to effectively capture the complex dynamic dependencies among observation points. In many real-world scenarios, the interactions among points in an irregular time series are not static; they evolve dynamically over time. Accurately capturing these dependencies is crucial for understanding the system’s behavior and making precise predictions (Chen et al., 2023; Zhang et al., 2022). However, existing methods generally rely on pre-defined and non-adaptive interaction structures. This means the scope of information exchange between observation points is often limited by a set of fixed, prior rules. For instance, ODE-based Interaction (Figure 2b), information interaction strictly follows the temporal order, meaning an observation can only influence its immediate subsequent state on the timeline and cannot establish a direct connection with more distant historical observations (Rubanova et al., 2019; Brouwer et al., 2019; Schirmer et al., 2022; Bilos et al., 2021). Meanwhile, graph-based methods (Figure 2c) construct connections based on fixed rules, such as belonging to the same time point or the same variable (Yalavarthi et al., 2024; Li et al., 2025a). The common limitation of these methods is their inability to dynamically and flexibly identify the truly relevant observation points based on the specific context of each point. As a result, they struggle to capture the deep dynamic correlations that span across time and variables.
@@ 2 Related Work
Recently, deep learning has made impressive progress in natural language processing (Liang et al., 2024; Li et al., 2025c; Zhou et al., 2024), computer vision (Lu et al., 2023, 2024a, 2024b; Zhou et al., 2025), and other aspects (Wang et al., 2025; Li et al., 2025b; Yang et al., 2024). Studies have shown that learned features may perform better than human-designed features. Therefore, many deep learning-based IMTSF methods have been proposed and have received substantial extensive attention. Currently, the related research can be broadly categorized into two main technical routes: structured representation-based methods and raw-data-based methods.
@@ 2.1 Structured Representation-based Methods
The core idea of these methods is to convert irregular, asynchronous data into a regular format through structural transformation, thereby making it compatible with standard sequence models. This category includes the following mainstream approaches: (1) Interpolation-based Methods: This technique generates new values at missing time points through function fitting to create an equally spaced time series, as adopted in works like mTAN (Shukla & Marlin, 2021). Its main limitation is the introduction of artificial data points that were not actually observed, which can alter the original data distribution and distort its intrinsic dynamic patterns (Shukla & Marlin, 2019; Zhang et al., 2023). (2) Time-Aligned-based Methods: This strategy maps the observations of all variables to a unified global timeline and fills in missing values, but this causes a loss of precise time interval information between original observations, thereby introducing distortion (Che et al., 2016; Rubanova et al., 2019; Brouwer et al., 2019). (3) Patch-Aligned-based Methods: To mitigate the sequence length problem caused by alignment, methods like t-PatchGNN (Zhang et al., 2024) divide the timeline into fixed-size patches for local alignment. However, this aggregation based on a predefined granularity may smooth out or lose critical fine-grained dynamics within each patch (Luo et al., 2025; Nie et al., 2023). In contrast, our ASTGI framework directly represents each discrete observation, fully preserving the original data and fundamentally avoiding the information distortion caused by structural transformations.
@@ 2.2 Raw-Data-based Methods
Unlike the previous category, this paradigm models the set of discrete observation points directly, avoiding the distortion introduced by data structuring. However, it typically relies on predefined, non-adaptive rules to capture the dependencies between points. (1) ODE-based Interaction: Represented by models like Latent-ODE (Rubanova et al., 2019) and NeuralFlows (Bilos et al., 2021), these methods treat the evolution of a time series as a continuous dynamical system. Although they can naturally handle queries at any time point, their inherent Markov assumption strictly confines interactions to temporally adjacent states. This prevents the model from capturing direct long-range dependencies between non-adjacent events (Brouwer et al., 2019; Schirmer et al., 2022). (2) Static Structure Interaction: These methods use Graph Neural Networks (GNNs) to learn the relationships between observation points. However, their graph structure is typically constructed based on fixed heuristic rules (Yalavarthi et al., 2024; Li et al., 2025a). Such a static topology is insensitive to the specific data context and cannot adapt as the system state evolves, making it difficult to capture event-driven dynamic associations. Unlike the fixed interaction rules of these methods, ASTGI adaptively constructs an interaction graph for each observation point, enabling it to dynamically capture context-dependent dependencies.
@@ 4.1 Experimental Setup
Datasets and Baselines. We conduct experiments on four widely-used public IMTS datasets: MIMIC, PhysioNet, Human Activity, and USHCN. To ensure fairness and comparability of the results, all datasets are preprocessed by strictly following the standard procedures established in prior state-of-the-art works (Yalavarthi et al., 2024; Zhang et al., 2024). We uniformly split the data into training, validation, and test sets with a ratio of 80%, 10%, and 10%, respectively. We selected a total of twelve state-of-the-art models from two main categories designed for irregular time series as baselines for comparison. For a detailed description of the datasets and baselines, please refer to the Appendix A.1 and A.2.
Implementation Details. All our experiments were conducted on a server equipped with an NVIDIA A800 GPU and implemented using the PyTorch 2.6.0+cu124 framework. All models are trained using the Mean Squared Error (MSE) as the loss function and optimized with the AdamW optimizer. We set the maximum number of training epochs to 300 and employ an early stopping strategy, where training is terminated if the model’s performance on the validation set does not improve for 5 consecutive epochs. To ensure a fair comparison across all models, we primarily adopted the hyperparameter settings reported in the original papers for the baseline models. Building on these configurations, we conducted further search and fine-tuning of key hyperparameters on the validation set for some models to ensure that each achieved a competitive level of performance. To ensure reproducibility and mitigate the effects of randomness, each experiment is run independently with five different random seeds (from 2024 to 2028), and we report the mean and standard deviation. Detailed hyperparameter configurations for all models are provided in the Appendix A.2.
@@ 5 Conclusion
Appendix A Detailed Experimental Setup
A.1 Datasets Details
@@
A.2 Baseline Model Details
@@
mTAN (Shukla & Marlin, 2021)

### v2

@@ 1 Introduction
Irregular Multivariate Time Series Forecasting (IMTSF) is a core problem across numerous critical scientific and engineering domains. Its applications are wide-ranging, from monitoring vital signs in intensive care units to tracking the evolution of environmental indicators in climate science (Yao et al., 2018; Vio et al., 2013; Shukla and Marlin, 2020; Brouwer et al., 2019; Bilos et al., 2021). This type of data is fundamentally defined by two characteristics: 1) intra-series irregularity, where observations of the same variable occur at unequal time intervals, and 2) inter-series asynchrony, where observation timestamps are misaligned across different variables. These traits make it challenging to directly apply existing models designed for regular time series. Furthermore, despite existing research dedicated to addressing irregular time series prediction tasks, current methodological paradigms are commonly constrained by two core challenges, which impede their application potential in complex real-world scenarios.
@@
The first challenge is how to accurately represent the original information of an irregular time series without introducing information distortion. An accurate representation must preserve the original sampling pattern, as this pattern itself contains key information about the system’s dynamics (Shukla and Marlin, 2020; Li and Marlin, 2020). Conversely, any artificial alteration to the original data structure introduces information distortion, corrupting this intrinsic information and ultimately compromising the model’s predictive performance (Ansari et al., 2023). There are three main representation paradigms, each with its own limitations. 1) Interpolation-based Methods (Figure 1b): This paradigm transforms an irregular series into an equally spaced series through numerical interpolation (Shukla and Marlin, 2021). However, this approach generates artificial data points that were not actually observed, which can introduce bias and distort the original sampling distribution (Shukla and Marlin, 2019; Zhang et al., 2023). 2) Time-aligned-based Methods (Figure 1c): This paradigm maps the observations of all variables to a unified timeline and fills in the missing values (Brouwer et al., 2019). Its main drawback is the loss of precise information about the time intervals between the original observations (Che et al., 2016; Rubanova et al., 2019). 3) Patch-aligned-based Methods (Figure 1d): This paradigm divides the time series into multiple patches (Nie et al., 2023). However, a rigid division may disrupt the continuity of information. Furthermore, intra-patch aggregation can smooth out critical, fine-grained dynamics (Zhang et al., 2024; Luo et al., 2025).
@@
The second challenge lies in how to effectively capture the complex dynamic dependencies among observation points. It is important to note that these two challenges are closely coupled in a progressive manner: accurate representation serves as the prerequisite and foundation for effective dependency modeling. If the raw information is distorted during representation (Challenge 1), the subsequent modeling of dynamic dependencies (Challenge 2) will inevitably be built upon inaccurate data, fundamentally compromising the model’s ability to capture true system dynamics. In many real-world scenarios, the interactions among points in an irregular time series are not static; they evolve dynamically over time. Accurately capturing these dependencies is crucial for understanding the system’s behavior and making precise predictions (Chen et al., 2023; Zhang et al., 2022). However, existing methods generally rely on pre-defined and non-adaptive interaction structures. This means the scope of information exchange between observation points is often limited by a set of fixed, prior rules. For instance, ODE-based Interaction (Figure 2b), information interaction strictly follows the temporal order, meaning an observation can only influence its immediate subsequent state on the timeline and cannot establish a direct connection with more distant historical observations (Rubanova et al., 2019; Brouwer et al., 2019; Schirmer et al., 2022; Bilos et al., 2021). Meanwhile, graph-based methods (Figure 2c) construct connections based on fixed rules, such as belonging to the same time point or the same variable (Yalavarthi et al., 2024; Li et al., 2025). The common limitation of these methods is their inability to dynamically and flexibly identify the truly relevant observation points based on the specific context of each point. As a result, they struggle to capture the deep dynamic correlations that span across time and variables.
@@ 2 Related Work
Existing methods for IMTSF can be broadly categorized into two paradigms based on how they handle data irregularity: Structured Representation-based methods and Raw-Data-based methods. The former transforms irregular data into regular structures to utilize standard sequence models, often at the cost of information distortion. The latter models discrete observations directly but typically relies on fixed rules for interaction. Our ASTGI framework falls into the second category but distinguishes itself by employing a fully adaptive graph interaction mechanism to capture dynamic dependencies without information loss.
@@ 2.1 Structured Representation-based Methods
The core idea of these methods is to convert irregular, asynchronous data into a regular format through structural transformation, thereby making it compatible with standard sequence models. This category includes the following mainstream approaches: (1) Interpolation-based Methods: This technique generates new values at missing time points through function fitting to create an equally spaced time series, as adopted in works like mTAN (Shukla and Marlin, 2021). Its main limitation is the introduction of artificial data points that were not actually observed, which can alter the original data distribution and distort its intrinsic dynamic patterns (Shukla and Marlin, 2019; Zhang et al., 2023). (2) Time-Aligned-based Methods: This strategy maps the observations of all variables to a unified global timeline and fills in missing values, but this causes a loss of precise time interval information between original observations, thereby introducing distortion (Che et al., 2016; Rubanova et al., 2019; Brouwer et al., 2019). (3) Patch-Aligned-based Methods: To mitigate the sequence length problem caused by alignment, methods like t-PatchGNN (Zhang et al., 2024) divide the timeline into fixed-size patches for local alignment. However, this aggregation based on a predefined granularity may smooth out or lose critical fine-grained dynamics within each patch (Luo et al., 2025; Nie et al., 2023). In contrast, our ASTGI framework directly represents each discrete observation, fully preserving the original data and fundamentally avoiding the information distortion caused by structural transformations.
@@ 2.2 Raw-Data-based Methods
Unlike the previous category, this paradigm models the set of discrete observation points directly, avoiding the distortion introduced by data structuring. However, it typically relies on predefined, non-adaptive rules to capture the dependencies between points. (1) ODE-based Interaction: Represented by models like Latent-ODE (Rubanova et al., 2019) and NeuralFlows (Bilos et al., 2021), these methods treat the evolution of a time series as a continuous dynamical system. Although they can naturally handle queries at any time point, their inherent Markov assumption strictly confines interactions to temporally adjacent states. This prevents the model from capturing direct long-range dependencies between non-adjacent events (Brouwer et al., 2019; Schirmer et al., 2022). (2) Static Structure Interaction: These methods use Graph Neural Networks (GNNs) to learn the relationships between observation points. However, their graph structure is typically constructed based on fixed heuristic rules (Yalavarthi et al., 2024; Li et al., 2025). Such a static topology is insensitive to the specific data context and cannot adapt as the system state evolves, making it difficult to capture event-driven dynamic associations. Unlike the fixed interaction rules of these methods, ASTGI adaptively constructs an interaction graph for each observation point, enabling it to dynamically capture context-dependent dependencies.
@@ 4.1 Experimental Setup
Datasets and Baselines. We conduct experiments on four widely-used public IMTS datasets: MIMIC, PhysioNet, Human Activity, and USHCN. To ensure fairness and comparability of the results, all datasets are preprocessed by strictly following the standard procedures established in prior state-of-the-art works (Yalavarthi et al., 2024; Zhang et al., 2024). We uniformly split the data into training, validation, and test sets with a ratio of 80%, 10%, and 10%, respectively. We selected a total of twelve state-of-the-art models from two main categories designed for irregular time series as baselines for comparison. For a detailed description of the datasets and baselines, please refer to the Appendix B.1 and B.2.
Implementation Details. All our experiments were conducted on a server equipped with an NVIDIA A800 GPU and implemented using the PyTorch 2.6.0+cu124 framework. All models are trained using the Mean Squared Error (MSE) as the loss function and optimized with the AdamW optimizer. We set the maximum number of training epochs to 300 and employ an early stopping strategy, where training is terminated if the model’s performance on the validation set does not improve for 5 consecutive epochs. To ensure a fair comparison across all models, we primarily adopted the hyperparameter settings reported in the original papers for the baseline models. Building on these configurations, we conducted further search and fine-tuning of key hyperparameters on the validation set for some models to ensure that each achieved a competitive level of performance. To ensure reproducibility and mitigate the effects of randomness, each experiment is run independently with five different random seeds (from 2024 to 2028), and we report the mean and standard deviation. Detailed hyperparameter configurations for all models are provided in the Appendix B.2.
@@ 5 Conclusion
Ethics statement
Our work exclusively uses publicly available benchmark datasets that contain no personally identifiable information. No human subjects are involved in this research.
Appendix A The Use of Large Language Models (LLMs)
We do not use Large Language Models in our methodology and writing.
Appendix B Detailed Experimental Setup
B.1 Datasets Details
@@
B.2 Baseline Model Details
@@
mTAN (Shukla and Marlin, 2021)
@@
Appendix C Qualitative Analysis of Learned Interaction Graphs
Figure 5: Visualization of the adaptively learned causal graph. The plot displays the interactions between observation points for a sample from the MIMIC dataset. The x-axis represents time (hours), and the y-axis represents different variables. Arrows indicate the direction of information flow (from history to query). The model successfully captures (1) synchronous correlations between variables (e.g., Bilirubin Direct and pH at $t=8$), (2) long-range temporal dependencies (e.g., Troponin I self-connection $t=4\rightarrow t=7$), and (3) cross-variable lagged effects. This confirms that ASTGI adaptively constructs a sparse and meaningful interaction topology.
To validate the interpretability and effectiveness of the proposed Neighborhood-Adaptive Graph Construction module, we visualize the learned causal graph for a randomly selected sample from the MIMIC test set.
Figure 5 illustrates the inference process. In this graph: (1) Nodes represent discrete observations, positioned horizontally by timestamp (Time) and vertically by variable type. (2) Node Size is proportional to the cumulative attention weight the node receives, indicating its importance in the current inference context. (3) Edges represent the learned attention scores. Darker and thicker lines indicate stronger dependencies. To ensure visual clarity, we only visualize the top-K strongest incoming edges for each node.
As shown in Figure 5, the learned structure exhibits three distinct patterns that align with the characteristics of irregular multivariate time series: (1) Synchronous Multivariate Correlations: The model frequently establishes strong connections between different variables observed at the same timestamp. For example, at $t=8$, a significant interaction is observed between Bilirubin Direct and pH (indicated by the thick connecting line). This demonstrates that ASTGI can effectively leverage synchronous co-occurrence information to reconstruct the system state, bypassing the need for manual time alignment. (2) Direct Long-Range Dependencies: Unlike recurrent models that propagate information step-by-step, ASTGI enables direct information propagation across non-adjacent timestamps. A clear example is the variable Troponin I, where the observation at $t=4$ strongly connects to the subsequent observation at $t=7$. This mechanism allows the model to retrieve critical historical information directly, thereby mitigating the long-term dependency issues common in recurrent architectures. (3) Cross-Variable Temporal Impact: The graph also captures complex lagged dependencies across different variables. We observe that the state of Troponin I at $t=4$ exerts a notable influence on Hemoglobin at $t=8$. This suggests that the model has learned to associate early anomalies in specific physiological indicators with delayed responses in others, capturing the systemic dynamics of the underlying process.
Appendix D Visualization of Learned Spatio-Temporal Embeddings
Figure 6: Visualization of the learned spatio-temporal embedding space. We visualize the coordinates $\bm{p}_{i}$ of observations from a MIMIC sample using t-SNE. Points are colored by variable types. The temporal evolution is indicated by the color intensity: for each variable, the color transitions from light (early time) to dark (late time). The clear clustering and continuous gradients demonstrate that ASTGI effectively encodes both variable semantics and dynamic temporal patterns.
To intuitively understand how ASTGI represents discrete irregular observations, we provide a visualization analysis of the learned spatio-temporal coordinate space. We randomly selected a test sample from the MIMIC dataset and extracted the coordinate vectors $\bm{p}_{i}=\bm{e}_{c_{i}}\oplus\bm{e}_{t_{i}}$ for all its observation points. We then utilized t-SNE to project these high-dimensional coordinates into a 2D space.
Figure 6 illustrates the resulting embedding structure. To clearly visualize the temporal evolution within the embedding space, we apply a time-dependent color gradient to the observation points of each variable. Specifically, lighter shades represent earlier observations, while darker and more saturated shades indicate later timestamps. The visualization highlights three key characteristics: (1) Variable Distinctiveness: Observations corresponding to different variables (represented by different colors) form distinct, well-separated clusters. This demonstrates that the learnable Channel Embedding effectively captures the unique semantic identities of different physiological indicators. (2) Temporal Continuity: For any given variable, the observation points do not collapse into a single spot but instead form a continuous trajectory. The smooth transition from light to dark colors confirms that our Time Encoding successfully preserves the sequential order and temporal intervals within the embedding space. (3) Validity for Adaptive Graph Construction: The combination of variable clustering and temporal trajectories creates a structured metric space. In this space, the Euclidean distance—used by our Neighborhood-Adaptive Graph Construction module—naturally prioritizes neighbors that are semantically correlated and temporally relevant. This validates our design choice of replacing fixed interaction graphs with dynamic k-NN search in this learned space.